Shift summary for Owl 12-10-99

Angela Bellavance (
Fri, 10 Dec 1999 05:31:10 -0600 (CST)

Shift Summary for owl shift, 10/Dec/99

On Shift: A. Bellavance

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7 (I assume)

Accounting of lost time: Tev still down

Runs ended this shift:none
Tapes brought to FCC this shift:none
Accesses during shift: D. Jensen to check on vacuum down process.
Overview of shift: walk-thrus
Description of Shift:

time comment
01:30 Angela arrives to find Sydney hard at work on something.
"Val" is complaining about gain drift - he is just silenced.
Red buttons on the SDAQ:
OnSplit - fine since J.Shields is doing his thing.
VolMon-CsI - Experts should probably double check this in the
morning. It complains that the MUX1 ENGTOTs are disabled.
What does that mean?
02:40 Doug checks up on the vacuum pump-down progress. He announces
that we are now allowed into the hall again.
03:00 Angela goes to the high-rise in search of a working color
05:00 Another walk-thru. Just me now. Most systems are fine.
SDAQ buttons mentioned above are still red. "Val" is
complaining again. There are many channels with gain drift
problems. The ones with largest drift are 799, 1096, and 2847.
Channels 2446 and 2447 look fine (note on board says to keep an
eye on them).

Follow up Notes on Previous Shift

We can access the hall again.

Items to Follow up on Next Shift

CsI people need to look at the gain drift.
Somebody needs to make sure the SDAQ VolMon-CsI red button isn't serious.

Angela Bellavance work: (630) 840-4144 home: (630) 548-3412 page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke